OpenStudy (jotopia34):

Another series question please: is the series: (sigma from n=2 to infinity)(-1)^n(4^-n) convergent or divergent, if convergent, what is the sum of the series

OpenStudy (jotopia34):

\[\sum_{n=2}^{\infty}(-1)^n\frac{ 1 }{ 4^n }\]

OpenStudy (jotopia34):

I see that n=2, which we usually don't have, we usually have n=0, but I don't know why you're supposed to change that.

OpenStudy (jotopia34):

\[\sum_{n=0}^{\infty}(-1)^{n+3}\frac{ 1 }{ 4^{n+2} }\]

OpenStudy (anonymous):

?

OpenStudy (jotopia34):

Now Im confused by the exponent "n" appearing outside the bracket and in the denominator

OpenStudy (anonymous):

You're making things too hard for yourself (take 2) \[\huge \sum_{n=2}^{\infty}(-1)^n\frac{ 1 }{ 4^n }=\sum_{n=2}^{\infty}\left(-\frac{ 1 }{ 4 }\right)^n\]

OpenStudy (jotopia34):

damn, overcomplication indeed

OpenStudy (anonymous):

I hope you have it from here :)

OpenStudy (jotopia34):

why would she tell us to take the n=2 down to zero. I hate when profs show you innecesary pellet

OpenStudy (anonymous):

I guess it's not that hard... \[\huge \sum_{n=2}^{\infty}(-1)^n\frac{ 1 }{ 4^n }=\sum_{n=0}^{\infty}\left(-\frac{ 1 }{ 4 }\right)^{n+2}\]

OpenStudy (anonymous):

Tell Professor Wendy I said hi ^_^

OpenStudy (jotopia34):

uh oh, I just realized I typed in the wrong equation. Its \[\sum_{n=2}^{\infty}(-1)^{n+1}/4^n\]

OpenStudy (anonymous):

As if that's an issue :P \[\huge \sum_{n=2}^\infty \frac{(-1)^{n+1}}{4^n}=(-1)\sum_{n=2}^\infty \frac{(-1)^{n}}{4^n}\]

OpenStudy (jotopia34):

ok, i see. youre negating the effect of the plus 1 in the exponent by taking out a neg 1 to the front

OpenStudy (anonymous):

So that the numerator and denominator have the same exponent :D

OpenStudy (jotopia34):

right, thats what I said ;)

OpenStudy (anonymous):

So, we're done?

OpenStudy (jotopia34):

ty!!!!!!

OpenStudy (jotopia34):

for now. I now thanks to you know how to finish this one

OpenStudy (anonymous):

Tell Professor Wendy she's getting too old... :D

OpenStudy (jotopia34):

will do.

OpenStudy (anonymous):

^_^